OpenStudy (anonymous):

A piece of tile is 4 1/2 inches long and 3 1/4 inches wide. What is the area of the tile?

OpenStudy (anonymous):

area = w* L

OpenStudy (anonymous):

=4 1/2*3 1/4=4.5*3.25

OpenStudy (anonymous):

=14.624

OpenStudy (anonymous):

14.625

OpenStudy (anonymous):

CHOSE FROM 7 3/4 OR 12SQU OR 12 1/8 OR 14 5/8

OpenStudy (anonymous):

ok 14 5/8

OpenStudy (anonymous):

0.625=5/8

OpenStudy (danjs):

\[\frac{ 9 }{ 2 } * \frac{ 13 }{ 4 } = \frac{ 117 }{ 8 } = 14 +\frac{ 5 }{ 8 }\]

OpenStudy (triciaal):

 unit of measure for area is square units